OpenStudy (anonymous):

calculate the roots of the eq x^2+x -1=0. I really dont want the roots but i want the quadratic to be expressed as multiplication of factors like (x- root1) (x-root2)

OpenStudy (anonymous):

Thanks for the answer which is quite hard to understand than using the simple \[(-b \pm \sqrt{b ^{2}-4ac})/2a \] when the quadratic is a\[ax ^{2}+bx+c=0\] but my doubt is: is the factor equation always expressed as (x-root1)(x-root2) or they do change the signs like (x-root1)(x+root2) something like that

OpenStudy (phi):

If you use the quadratic formula you get two roots, called r1 and r2 then your factors would be (x-r1)(x-r2)

OpenStudy (asnaseer):

@phi is right. the reason you can also end up with something like (x-a1)(x+a2) is if one of the roots is a negative number. e.g., 1. say the roots were 3 and 7, then this would give you (x-3)(x-7) = 0 2. say the roots were -3 and 7, then this would give you (X+3)(x-7) = 0 3. etc...

OpenStudy (asnaseer):

another way of writing my example number 2 is: (x-(-3))(x-7) = 0

OpenStudy (jamesj):

If r1 and r2 are roots of ax^2 + bx + c = 0, then it must be the case that ax^2 + bx + c = a(x - r1)(x-r2) = 0

OpenStudy (asnaseer):

yes, I agree with @JamesJ, I was trying to address @chaotic.clone's question about whether you could end up with (x-something)(x+somethingelse). I don't know if I was clear enough.

OpenStudy (anonymous):

@JamesJ, is it true that if the eq is ax^2+bx+c=0 has the roots of form just (x-r1)(x-r2) but not a(x-r1)(x-r2) as you have said.

OpenStudy (jamesj):

What I am saying is that as polynomials, ax^2 + bx + c = a(x - r1)(x-r2). The a is strictly necessary, even if it is usually that case that a = 1.